How much time will it take to gain $270 in simple interest with an interest rate of 5% and an investment of $1,200?
Answer by pakhi(24)

The equation for simple interest calculation is
S.I. = (P * R * T)/100
where S.I. = simple interest
P = principal amount
R = rate of interest in percentage
T = time in years
'*' symbol denotes multiplication

Therefore according to the problem
270 = (1200 * 5 * T)/100
or 270 * 100 = 6000T
or T = 27000/6000 = 4.5yrs

So the required time to get $270 as simple interest on $1200 @ 5%is 4years and 6months.